Previous 1 2 34567 8 9 10 Question 7 of 27 (1 point) 5.3 Section Exercise 35 A fair die is rolled six times. What is the probability that it comes up 6 at least once? Round the answer to four decimal places. The probability that it comes up 6 at least once is Gr As Li
Answer #1

Solution:

To find out probability that you will get a 6 at least once, we need to find out the probability that you will NOT get a 6 any of those times

p = probability of getting a 6 on each throw = 1/6.

p' = probability of NOT getting a 6 on each throw is  1-p = 5/6.

The probability of not getting a 6 n times = p' to the nth power.

= (5/6) ^6

= 15,625 / 46,656

=0.334

The probability of 6 getting atleast once = 1- 0.334 =0.666